### v1

# Practical aspects for the creation of an audio dataset from field recordings with optimized labeling budget with AI-assisted strategy

Abstract
Machine Listening focuses on developing technologies to extract relevant information from audio signals. A critical aspect of these projects is the acquisition and labeling of contextualized data, which is inherently complex and requires specific resources and strategies. Despite the availability of some audio datasets, many are unsuitable for commercial applications. The paper emphasizes the importance of Active Learning (AL) using expert labelers over crowdsourcing, which often lacks detailed insights into dataset structures. AL is an iterative process combining human labelers and AI models to optimize the labeling budget by intelligently selecting samples for human review. This approach addresses the challenge of handling large, constantly growing datasets that exceed available computational resources and memory. The paper presents a comprehensive data-centric framework for Machine Listening projects, detailing the configuration of recording nodes, database structure, and labeling budget optimization in resource-constrained scenarios. Applied to an industrial port in Valencia, Spain, the framework successfully labeled 6540 ten-second audio samples over five months with a small team, demonstrating its effectiveness and adaptability to various resource availability situations.
Keywords: Machine Listening, Active Learning, Artificial Intelligence, Dataset, IoT
1 Introduction
Machine Listening is the field of knowledge that aims to create technological solutions that are able to obtain relevant information from audio signals. Some of the most common problems present in this field can be Acoustic Scene Classification (Schmid et al. (2023)), Sound Event Detection (Mesaros et al. (2021); Khandelwal et al. (2024)) or Anomalous Sound Detection (Wilkinghoff (2024)) among others (Aguado et al. (2023)).
One of the main pillars of any Machine Listening project, and of any other AI project, is the acquisition and labeling of relevant and contextualized data for the problem to be addressed. This task is usually not simple and needs specific resources and strategies to be carried out successfully. There are several projects in the literature related to the creation of generic sound databases to be used in Machine Listening projects. The Sounds Of New York (SONYC) (Mydlarz et al. (2019)) project is one of them, in which about 50 recording nodes were deployed throughout various areas of New York City. Other similar projects, in which nodes have also been deployed in areas of high urban concentration, are those of Montevideo (Zinemanas et al. (2019)) and Singapore (Ooi et al. (2021)).
Although there are some audio datasets, they are not usually suitable for commercial purposes. The most widely used in the scientific literature is Audioset (Gemmeke et al. (2017)) which is composed of audios from YouTube videos. This dataset, while extensive, does not give access to the audios themselves but only to the feature vectors. Furthermore, it can only be used for research activities, never for commercial uses and purposes. Even though some of them can be used for commercial purposes, no updates and/or improvements to these datasets have been released for a long time (Piczak (2015); Salamon et al. (2014)).
As for labeling, two common strategies are observed: Active Learning (AL) based on expert opinions or crowdsourcing (Martín-Morató and Mesaros (2023); Lipping et al. (2019); Martín-Morató et al. (2021)). In the case of crowdsourcing, the usual scenario is that the labelers are not experts in the topic, so that no real insight into the structure and details of the dataset is obtained. In the opinion of the authors of this paper, the best option is to work through AL iterations with a pool of expert human labelers.
Since the usual scenario is that the amount of data to be labeled far exceeds the resources available to perform the task, the procedure known as Active Learning aims at optimizing the labeling budget. Active Learning is an iterative procedure that combines human labelers and AI models in such a way that these models smartly select which samples should be listened and labeled by humans. This selection is done by optimizing a certain criterion based on the principles of uncertainty, discrepancy of a model committee, estimation of change in the model, among others (Settles (2009)). In published scientific work on AL for audio, it is common for the dataset to have a controlled and computationally manageable size, both from the point of view of computational complexity and memory space. In our case study this is not the case because the dataset grows constantly so that there comes a point when both the computational resources and the memory available are not enough to cope with the necessary calculations on the entire dataset. In this paper a series of measures are proposed and implemented in order to deal with this situation.
This paper presents a complete framework to approach Machine Listening projects from a Data-Centric perspective (see Figure 1(a)). The proposed work presents the configuration of the recording node, the structure of the database that compose the dataset and the optimization of the labeling budget in a scenario of constrained resources (small labelers team) when it is intended to avoid crowdsourcing. The presented framework can be parameterized to be modified depending on the particular case and availability of resources. The results shown in this work are obtained from the use case applied to the context of an industrial commercial port located in the city of Valencia (Spain) where 7 recording nodes are currently deployed. At the time of writing this paper, 6540 audios of 10 seconds duration have been labeled during 5 months (since January 8, 2024) with an effort of 30 minutes of labeling per day, which corresponds to each tagger labeling 40 audios per day. The human labelers teams is composed of five people which have been divided into two groups in order to perform cross-validation techniques and avoid biases.
Figure 1: (a) Steps to be taken to solve Machine Listening problems with a Data-Centric perspective. (b) Recording node block diagram. (c) Recording node real ensemble
2 Materials and Methods
The framework is formed by: the recording node, the raw storage of the audio data, the AL process and the building of a structured database with the labeled data. The following sections present each of the modules.
2.1 Recording node
The core of the IoT device for recording is a Raspberry Pi card. Additional peripheral devices were added to improve remote monitoring and avoid possible failures in the event of a power outage (see Figures 1(b) and 1(c)): (i) a Rode Lavalier II microphone, (ii) a USB sound card to connect the microphone to a USB slot, (iii) an SPS Hat so that a safe shutdown occurs in the event of power line supply failure and (iv) a USB router with internet connectivity through a SIM card to remotely monitor the node’s activity. In the case of power supply through solar panels, the SPS Hat is indispensable as it allows setting hardware on and off times. Once assembled, the device is placed inside a PVC IP65 box with two openings: one for the microphone and the other for the power supply cable.
When recording, files are saved every 10 seconds. The timestamp of the 10-second chunk is implicit in its filename. Remote monitoring of the nodes was implemented via the periodic transmission of different node state variables using the MQTT protocol. A ThingsBoard panel was implemented to visualize these states. To detect possible failures of the microphone, a periodic sending of an audio file every hour has been scheduled to a Minio bucket. Thus, every day, a specific number of audios from each node are listened to verify the microphone status. For illustrative purposes, a repository with the described functionalities has been released [^1] https://github.com/JNaranjo-Alcazar/recording_node_monitoring.git.
2.2 Active Learning
Several AL solutions have been proposed in recent years in the context of Machine Listening solutions. The line of works that concludes with Shuyang et al. (2020) can be understood as an evolution of an AL algorithm intended for audio classification. In Shuyang et al. (2017) an AL method called Medoid Based AL (MAL) was proposed based on performing an unsupervised K-medoid clustering on unlabeled data. The medoids are then proposed to be labeled by human annotators. In each iteration, the medoids are excluded from the next proposal. The next work, called MAL-MF (Mismatch-First) (Shuyang et al. (2018)), can be considered an improvement of the MAL method. In MAL-MF, MAL is used for the first iteration, when there is no human-labeled data. Once labeled data is available, a classifier is trained. From the second iteration onwards, a committee-based approach is used where samples with discrepancies between the classifier and the medoids propagation are candidates for labeling. Samples with discrepancies are sorted according to a distance criterion for diversity consideration.
Other approaches are based on the uncertainty of a certain classifier (Settles (2009)). The main idea of uncertainty-based approaches is that the unlabeled samples with the lowest classification probability to a specific class should be labeled next. The classifier is trained iteratively as new labeled samples are available. Works that employ this method include using Monte Carlo dropout (Shishkin et al. (2021)), Gaussian distribution for estimated informativeness (Shishkin et al. (2024)) or Meire et al. (2023) which studies, in addition to uncertainty, a hybrid approach using semi-supervised detection of outliers techniques. Although uncertainty-based approximations provide very informative samples, these approximations can lead to bias due to mislabeling. To avoid this phenomenon, in Wang et al. (2019) a solution is proposed to label certain samples every a specified number of iterations.
In this work the MAL-MF was implemented. However, uncertainty principles with a pre-trained network are used to optimize the AL procedure when a large amount of data is available. Information from each AL iteration is stored in an intermediate database.
2.2.1 Active Learning Database
The tables that compose this database are:
  • ALPreprocessing: information about the time window for analysis, the number of audios, the number of folds, the date and the percentage of labeled data that contains the execution (see step 1 in Fig.2).
  • WavsProposed: information relevant to the audios proposed by the AL algorithm. It contains two foreign keys: (i) the ID of the AL iteration (id of ALPreprocessing) and (ii) the id of the proposed audio from the table Audios. Other fields, such as the label, the number of labelers who have tagged the audio and the percentage of agreement of the labelers to tag the audio, the file name and the node’s id are also included.
2.3 Audio Database
This database is composed of the information obtained in the labeling process. It is constituted by the following tables:
  • Nodes: information associated to each deployed IoT Recording Node.
  • Paths: paths of the folders where the audios are stored.
  • Ontology: tags present in the database
  • Labelers: Id of the human labelers
  • Audios: information about the audios captured by the recording nodes such us filename, sampling rate, bits per sample, duration, or number of channels. It has two foreign keys: (i) the path’s and (ii) node’s id.
  • Chunks: all audio events identified in the labeling process. It contains three foreign keys: (i) id of the corresponding audio file, (ii) id of the event label from the Ontology table and (iii) the labeler’s id. In the case of labeling the beginning and end of the event (strong labeling), two fields corresponding to the onset/offset times of the event must be added.
Figure 2: Framework for creating an audio database with a large amount of available data while optimizing the labeling budget
2.4 Detailed Active Learning Framework
Our AL implementation follows the one proposed in Shuyang et al. (2018). Many of the AL algorithms in the literature are suitable only for problems with small-scale datasets where all the samples can be used for calculations in each iteration of the AL procedure. However, the presented scenario faces a huge and constantly growing unlabeled dataset, thus, there will come a point where the algorithm cannot be applied to the entire data set due to computational capacity limitations. Therefore, measures must be taken to deal with this situation. One possibility is to perform the AL iterations over time windows, e.g., a particular month, a particular week, etc. (step 1 in Figure 2). Even with this restriction the amount of data can still be very large. As a solution to this, the procedure proposed, and implemented, in this paper is based on dividing the samples into different disjoint sets. The criterion to assemble the sets is based on the variability shown in the CNN14 PANNs top-1 predictions (the class with the highest probability), where PANNs is a collection of models of different CNNs with public checkpoints trained with Audioset (Kong et al. (2020)). The sets are located in different folders (one folder per set) in such a way that the first folder is more informative and diverse than the last one. This division process is depicted in steps 3 and 4 of Figure 2 and is explained in more detail below. The available computational capacity constrains the quantity of samples in each folder.
The first classifier of the committee corresponds to the propagation of labels from human-labeled samples, which are considered the medoids of a given cluster, to their nearest unlabeled neighbors. The second classifier is a Logistic Regression. The feature vector used to train both, the Logistic Regression and the medoid propagation, is the PANNs pre-trained network CNN14 embedding. An uncertainty approach is used for the division of samples among folders. The steps performed for an AL iteration (see Figure 2) are:
  1. 1.
    Due to the large amount of available data recorded by the recording nodes the AL framework should be performed over a finite and defined time window.
  2. 2.
    Of the audios selected in step 1, those that have already been tagged by humans are used as medoids.
  3. 3.
    The PANNs network is used to obtain a first classification of the samples under study. Only the label with the highest probability is assigned to that sample.
  4. 4.
    As the number of samples may still be too large to be processed by the available computational capacity some further partition of the data may be necessary. Samples are disjointly distributed in different sets (folders). The samples are distributed in the different sets in such a way that they (the sets) can be ordered from greater to lesser diversity and uncertainty. Diversity means that, if very few samples of a given class have been found, they are concentrated in the first sets. That is, the sets that should be treated first. By uncertainty, we mean that if there are many samples of a given class and these samples are distributed in different sets, the ones with the lowest probability, according to the PANNs classification, will be placed in the first folders.
  5. 5.
    The number of medoids can also be a limiting factor for the computational capacity. Thus, a selection criterion must be used. In our implementation, if there are labeled samples from the same recording node and from the same time window, they are added as medoids first. In case more medoids can be added, labeled samples from the same node but from different windows are used. Finally, samples from other nodes already labeled would be added. After the selection of medoids the AL algorithm is launched on the selected folder (Shuyang et al. (2018)).
  6. 6.
    AL database tables are updated
  7. 7.
    The samples proposed by the AL algorithm are provided to the human-labeler teams. BAT [^2] https://github.com/BlaiMelendezCatalan/BAT has been used as the labeling tool in this work.
  8. 8.
    Chunks table of the database is updated as well as the medoids field of the ProposedWavs table of the AL database. Since several tags may appear in a single audio, a strategy must be defined to define which is the medoid class of the audio for propagation in the next iteration of AL. The criterion followed for updating the medoids field with a label depends on the agreement between the labelers within its group. This agreement must satisfy two requirements: (i) the labels have to be in, at least, 2/3 of the labelers’ annotations within the group and (ii) if more than one label satisfies this requirement, the selected label will be the one that lasts the longest.
3 Result
The labeling team was composed of 5 people divided into two groups of 3 and 2 members. In each AL iteration, 400 audios were proposed. Each group tagged 200 audios per week, 40 audios per day. Thus, the daily listening time corresponded to about 7 minutes per labeler, estimating a labeling effort of 30 minutes per day. Every Friday, a 30-minute meeting was held to unify criteria, propose suggestions and resolve doubts among taggers. The ontology used as a starting point was the Audioset ontology (Gemmeke et al. (2017)). If a labeler considered that an audio event could not be labeled with the available classes, a suggestion method was implemented so that the class became available in the next labeling iteration. These Friday meetings were also used to notify the rest of the labelers of the incorporation of new classes into the labeling process. In addition, a Doubt class was created to be used in case a labeler was unsure which class to assign to a specific event. Every 10 labeling iterations, a doubt resolution iteration is performed. This iteration consisted in that each labeler had to re-listen and re-label their audios with some Doubt event from the previous iterations and resolve them. The tagging process started on January 8, 2024. When this paper was drafted, 18 AL iterations and 1 doubt resolution iteration have been performed. The results obtained correspond to a total of 6540 strong labeled audios (due to vacations, not every week 400 audios could be labeled) which corresponds to a duration of 18 hours of labeled audio having employed 171 different labels. The tagging effort corresponds to 52 hours per labeler. It should be noted that the creation of the groups can be parameterized and even (if deemed appropriate) different audios could be proposed to each labeler avoiding cross-validation, thereby increasing the number of labeled audios with the equivalent effort.
Visual inspection using the UMAP dimensionality reduction algorithm has been used to validate the correct operation of the AL cycle. A concrete iteration of the AL process is shown in Figure 3(a). The medoids used in that iteration, the audios proposed for labeling and the discarded audios are plotted. In this particular iteration, the 10th, 400 audios have been proposed. As can be seen in the figure, the proposed audios are spread all over the feature space, some of them being close medoids. To analyze if this proposition makes sense, the same feature space is plotted in Figure 3(b) but indicating the class provided by the CNN14 of PANNs pre-trained with Audioset. As can be appreciated (circled on the image) the AL solution proposes the border samples between two classes. In addition, audios that are classified as Silence (green dots) have not been proposed in this iteration since medoids are available and they are not conflicting samples. By visually analyzing this iteration, it can be observed that the AL procedure is proposing relevant samples.
Figure 3: (a) 2D UMAP representation of an AL iteration. The orange dots represent the medoids used in the iteration, the green stars correspond to the audios proposed for labeling and the remaining dots are the audios discarded for labeling. (b) 2D UMAP representation top1 pre-trained PANNs prediction. The different colors show distinct classes. The circles mark areas of interest according to the AL process.
4 Conclusion
This paper establishes a complete framework for the creation of an audio database while optimizing the labeling budget in high-volume unlabeled data environments. This framework could be extremely useful when crowdsourcing is to be avoided to keep insight during the labeling process. The use case presented has been carried out in a commercial-industrial environment such as the port of Valencia (Spain). At the time of drafting this paper, 6540 10-second audios have been tagged with an effort of 52 hours per labeler having a pool of 5 labelers divided into two groups of 2 and 3 members. A dataset is expected to be released soon as one of the outcomes of this work.
Broader Impact Statement
The work proposes a framework to be used for Machine Listening projects in scenarios with constrained tagging budgets. General aspects of data collection, the AL-based labeling with AI-assisted techniques and the storage of the information obtained from the labeling process are presented. In Machine Listening projects it is common that the largest and best-known sounding databases are either not suited to the particular problems of the project or cannot be used for commercial purposes. That is why a framework such as the one presented here should become standard practice in any AI-based machine listening project.
Acknowledgments and Disclosure of Funding
The participation of Javier Naranjo-Alcazar, Jordi Grau-Haro and Pedro Zuccarello in this research was funded by the Valencian Institute for Business Competitiveness (IVACE) and the FEDER funds by means of project Soroll-IA2 (IMDEEA/2023/91).